# 979-8-3503-9086-5/24/\$31.00 ©2024 IEEE

### v1

@@ II-A Wavelet transform
where: $n$ is the discrete time index, $x[n]$ is the discrete signal of length $N$. CWT for a time-discrete signal is computed by the discrete summation of the dot product within the sampling interval. The translation parameter b and scale parameter a are in continuous forms, where translation is sample-wise, and scale spans a range of continuous natural numbers. This process produces a coefficient matrix of size (N, a), where N is the data length and a is the scale. The scalogram of CWT is illustrated in Fig. 1(a), with the vertical direction representing the frequency/scale ($\omega$/a) and the horizontal direction representing the time/translation (t/b).
The CWT is commonly implemented using the PyWavelets library, specifically through the pywt.cwt function. However, this function does not provide options for adjusting the wavelet kernel length or customizing the scalogram size based on specific requirements. This limitation results in a high computational cost, particularly for tasks that require low temporal resolution. Therefore, an approach that optimizes computational efficiency while preserving essential time-frequency features is highly desirable.
@@ II-B Proposed idea
Secondly, rather than producing a full output length equivalent to the input signal for each scale in the CWT, the proposed approach samples the output using a predefined hop size (H). This strategy reduces the number of extracted temporal features while lowering computational complexity during the CWT process. By performing a grid search over the parameters WL and H, users can optimize computational cost at satisfactory performance in trained models.
@@ III-A Dataset
The present study uses the MIMII dataset [37], which comprises acoustic recordings collected from industrial environments. The dataset includes sound data from various types of machinery, specifically fans, pumps, sliders, and valves. For each machine type, the dataset categorizes audio recordings into two classes: normal, representing properly functioning machinery, and abnormal, corresponding to malfunctioning machinery. The primary objective of utilizing this dataset is to develop and train models for the detection of machinery faults based on audio signals. To enhance robustness, the recordings are superimposed with environmental background noise at three distinct levels of signal-to-noise ratio (SNR): 6 dB, 0 dB, and 6 dB. The dataset consists of a total of 54,507 audio files, each with a duration of 10 seconds. All recordings are sampled at 16 kHz, yielding 160,000 samples per audio file.
@@ III-B Scalogram generation
The process of scalogram generation is illustrated in Fig. 2, and is herein referred to as optCWT (optimal continuous wavelet transform). Initially, a wavelet kernel of length WL is defined at a given scale S. By specifying the wavelet length, the transformation can be tailored to balance the trade-off between prediction performance and computational cost. Specifically, longer wavelets may enhance model performance at the expense of increased computational complexity, whereas shorter wavelets may reduce computational demands with potentially lower performance. Subsequently, the discrete audio signal x(n) of length N, as obtained from the aforementioned dataset, is convolved with the wavelet kernel using the Fast Fourier Transform (FFT), in accordance with the standard wavelet transform procedure. The resulting output is then downsampled by a factor of H, rather than retaining its original resolution. This downsampling operation reduces both memory usage and computational overhead, thereby optimizing the efficiency of subsequent processing steps, such as heat maps visualization. The transformation is iteratively performed for the same audio signal across the specified range of wavelet scales. Upon completion, a coefficient matrix of dimensions (N/H, S) is produced, which is more compact compared to the standard wavelet transform coefficient matrix of size (N, S). This reduced matrix is subsequently utilized to generate heatmaps, referred to as scalograms. The implementation for scalogram generation has been made publicly available via the GitHub repository referenced in [38]. For visual comparison, the scalograms produced by both the conventional CWT and the proposed optCWT are presented side by side in Fig. 1 (a) and (b). As observed, the CWT-based scalogram exhibits a more continuous and densely concentrated energy distribution, providing a more detailed representation. In contrast, the optCWT scalogram displays a sparser and more discrete distribution of energy, reflecting the trade-offs introduced by the optimized transformation.
@@ III-C Acoustic recognition task
To perform the acoustic recognition task, a CNNs model is developed and use generated scalograms images in above step as input data. The model is trained and tested for performance of fault detection via acoustic recognition. The performance of the optCWT-generated scalograms on the CNNs is evaluated and compared to that of scalograms produced by the conventional CWT to assess the efficacy of optCWT in detecting anomalous sounds. The developed pipeline is identical to the benchmark study [2] for comparison purpose. The setting for scales, heat maps generation, parameters of CNNs is also referenced. The only extra steps is the definition of wavelet kernel length and sampling the output of CWT. Two main parameters of this study, wavelet length WL and hop size H are examined in grid search manner to find an optimal set WL = 64, H = 128, which is significantly reduce the computation cost while achieving a satisfactory performance.
The research employs the PyWavelets library [39] for wavelet transformation, the TensorFlow library for implementing the binary classification CNNs model, and the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) [40] as the prediction performance metric.
@@ III-D Results
Overall, optCWT delivers performance that is fair and comparable to that of the baseline model, indicating its potential applicability in practical scenarios. Moreover, optCWT offers a substantial computational advantage as recorded in Table V and illustrated in Fig. 7. In the experimental setup, the generation of optCWT for a single file required only 1.15 seconds, whereas the baseline’s CWT required 8.09 seconds - approximately seven times longer. For the complete dataset comprising 54,507 files, the total processing time amounted to 17.41 hours for optCWT, compared to 121.5 hours for the baseline CWT. These results underscore the significant computational efficiency of optCWT.

### v2

@@ II-A Wavelet transform
where: $n$ is the discrete time index, $x[n]$ is the discrete signal of length $N$. CWT for a time-discrete signal is computed by the discrete summation of the dot product within the sampling interval. The translation parameter b and scale parameter a are in continuous forms, where translation is sample-wise, and scale spans a range of continuous natural numbers. This process produces a coefficient matrix of size (N, a), where N is the data length and a is the scales range. The scalogram of CWT is illustrated in Fig. 1(a), with the vertical direction representing the frequency/scale ($\omega$/a) and the horizontal direction representing the time/translation (t/b).
The CWT is commonly implemented using the PyWavelets [37] library, specifically through the pywt.cwt function. However, this function does not provide options for adjusting the wavelet kernel length or customizing the scalogram size based on specific requirements. This limitation results in a high computational cost, particularly for tasks that require low temporal resolution. Therefore, an approach that optimizes computational efficiency while preserving essential time-frequency features is highly desirable.
@@ II-B Proposed idea
Secondly, rather than producing a full output length equivalent to the input signal for each scale in the CWT, the proposed approach samples the output using a predefined hop size (H) in convolution step of each scale. This strategy reduces the number of extracted temporal features while lowering computational complexity during the CWT process. By performing a grid search over the parameters (WL) and (H), users can optimize computational cost at satisfactory performance in trained models.
@@ III-A Dataset
The present study uses the MIMII dataset [38], which comprises acoustic recordings collected from industrial environments. The dataset includes sound data from various types of machinery, specifically fans, pumps, sliders, and valves. For each machine type, the dataset categorizes audio recordings into two classes: normal, representing properly functioning machinery, and abnormal, corresponding to malfunctioning machinery. The primary objective of utilizing this dataset is to develop and train models for the detection of machinery faults based on audio signals. To enhance robustness, the recordings are superimposed with environmental background noise at three distinct levels of signal-to-noise ratio (SNR): -6 dB, 0 dB, and 6 dB. The dataset consists of 54,507 audio files, each with a duration of 10 seconds. All recordings are sampled at 16 kHz, yielding 160,000 samples per audio file.
@@ III-B Scalogram generation
The process of scalogram generation is illustrated in Fig. 2, and is herein referred to as optCWT (optimal continuous wavelet transform). Initially, a wavelet kernel of length WL is defined at a given scale S. By specifying the wavelet length, the transformation can be tailored to balance the trade-off between prediction performance and computational cost. Specifically, longer wavelets may enhance model performance at the expense of increased computational complexity, whereas shorter wavelets may reduce computational demands with potentially lower performance. Subsequently, the discrete audio signal x(n) of length N, as obtained from the aforementioned dataset, is convolved with the wavelet kernel using the Fast Fourier Transform (FFT), in accordance with the standard wavelet transform procedure. The resulting output is then downsampled by a factor of H, rather than retaining its original resolution. This downsampling operation reduces both memory usage and computational overhead, thereby optimizing the efficiency of subsequent processing steps, such as heat maps visualization. The transformation is iteratively performed for the same audio signal across the specified range of wavelet scales. Upon completion, a coefficient matrix of dimensions (N/H, S) is produced, which is more compact compared to the standard wavelet transform coefficient matrix of size (N, S). This reduced matrix is subsequently utilized to generate heatmaps, referred to as scalograms. The implementation for scalogram generation in this study has been made publicly available via the GitHub repository referenced in [39], and a pull request proposing the findings to upgrade the Pywavelets library is already created [40]. For visual comparison, the scalograms produced by both the conventional CWT and the proposed optCWT are presented side by side in Fig. 1 (a) and (b) respectively. As observed, the CWT-based scalogram exhibits a more continuous and densely concentrated energy distribution, providing a more detailed representation. In contrast, the optCWT scalogram displays a sparser and more discrete distribution of energy, reflecting the trade-offs introduced by the optimized transformation.
@@ III-C Acoustic recognition task
To perform the acoustic recognition task, a CNNs model is developed and uses the scalogram images generated in the previous step as input data. The model is trained and tested for performance of fault detection via acoustic recognition. The performance of the optCWT-generated scalograms on the CNNs is evaluated and compared to that of scalograms produced by the conventional CWT to assess the efficacy of optCWT in detecting anomalous sounds. The developed pipeline is identical to the benchmark study [2] for comparison purpose. The setting for scales, heat maps generation, parameters of CNNs is also referenced. The only extra steps are the definition of wavelet kernel length and sampling the output of CWT. Two main parameters of this study, wavelet length WL and hop size H are examined in grid search manner to find an optimal set WL = 64, H = 128, which is significantly reduce the computation cost while achieving a satisfactory performance.
The research employs the PyWavelets library for wavelet transformation, the TensorFlow library for implementing the binary classification CNNs model, and the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) [41] as the prediction performance metric.
@@ III-D Results
Overall, optCWT delivers performance that is fair and comparable to that of the baseline model, indicating its potential applicability in practical scenarios. Moreover, optCWT offers a substantial computational advantage as recorded in Table V and illustrated in Fig. 7. In the experimental setup, the generation of optCWT for a single file required only 1.15 seconds, whereas the baseline’s CWT required 8.09 seconds - approximately seven times longer. For the complete dataset comprising 54,507 files, the total processing time amounted to 17.5 hours for optCWT, compared to 122.5 hours for the baseline CWT. These results underscore the significant computational efficiency of optCWT.